# End-to-End Speech-to-Text Translation: A Survey

### v1

@@ 2.3 Loss Functions
MuST-C
En
14 lang
0.4K
1.6K
Librispeech
En
Fr
0.2K
1.4K
CoVost
En
11 lang
0.7K
11K
CoVost2
21 lang
En
2.8K
11K
En
15 lang
0.7K
78K
EuroparlST
4 lang
4 lang
0.25K
VoxPopuli
En
15 lang
1.79K
4.3K
Kosp2e
Ko
En
0.2K
0.2K
GigaST
En
De, Zh
10K
Prabhupadavani
en-bn-sn code-mix
25 lang
0.09K
0.13K
How2
En
Pt
2K
3 Datasets for ST Tasks
@@
    MuST-C (Di Gangi et al., 2019) It is a large multilingual ST translation corpus available . It contains translations from English into fourteen additional languages and is compiled from TED Talks. mTEDx (Salesky et al., 2021) is one such multilingual dataset from TED talks.
@@
    Kosp2e (Cho et al., 2021a) is a Korean (ko) to English(en) ST translation corpus, which contains Korean speech with parallel English texts. The corpus contains data from four different domains: Zeroth from news/newspaper, KSS (Park, 2018) from textbooks, emphStyleKQC (Cho et al., 2021b) from AI applications, and Covid-ED (Lee et al., 2021) from covid diaries of people which have emotions.
@@
Besides these popular ST datasets, there are some other smaller size datasets such as Fisher(Cieri et al., 2004), Call-Home [^4] https://ca.talkbank.org/access/CallHome/eng.html, Gordard Corpus(Godard et al., 2017), Glosse Audio Corpus [^5] https://ainu.ninjal.ac.jp/folklore/en/, BTEC  [^6] http://universal.elra.info/product_info.php?cPath=37_39&products_id=80, WSJ [^7] https://catalog.ldc.upenn.edu/LDC93s6a, IWSLT [^8] https://iwslt.org/, CHiME-4 Corpus(Christensen et al., 2010), Miami Corpus(Deuchar, 2008), and MSLT Corpus (Federmann and Lewis, 2016).
4 Toolkits for ST
In order to facilitate building and training ST models, various researchers have proposed a few toolkits. The toolkits for ST create an environment where the dataset for ST tasks can be pre-processed, and models can be trained, fine-tuned, and evaluated. We provide a short description of these toolkits to make the survey a place for a one-stop-shop for ST modeling.
@@ 4 Toolkits for ST
    EspNet-ST [^9] https://github.com/espnet/espnet toolkit(Inaguma et al., 2020) is developed as there was no toolkit available for performing the sub-tasks of ST. EspNet-ST provides ASR, LM, E2E-ST, Cascade-ST, MT, and TTS along with examples. It also provided pre-trained transformer-based models on various datasets like MUST-C, Libri-trans, Fisher, CALL-HOME, and How2.
@@
  • 3.
    NeurST [^11] https://github.com/bytedance/neurst (Zhao et al., 2020) is a lightweight toolkit, as it has no dependency on kaldi[CITE]. It has high computation efficiency using mixed precision and accelerated linear algebra and achieves faster training on large-scale datasets using Horovod and Byteps [CITE].

### v2

@@ 8 Deployment of E2E ST Models
MuST-C En 14 lang 0.4K 1.6K
Librispeech En Fr 0.2K 1.4K
CoVost En 11 lang 0.7K 11K
CoVost2 21 lang En 2.8K 11K
En 15 lang 0.7K 78K
EuroparlST 4 lang 4 lang 0.25K
VoxPopuli En 15 lang 1.79K 4.3K
Kosp2e Ko En 0.2K 0.2K
GigaST En De, Zh 10K
Prabhupadavani en-bn-sn code-mix 25 lang 0.09K 0.13K
How2 En Pt 2K
FLEURS 102 lang 102 lang 1.4K 0.3K
BSTC Zn En 98
Indic-TEDST En 9 lang 189 1.64K
9 Resources for ST
9.1 Datasets for ST Tasks
@@
    MuST-C (Cattoni et al., 2021) It is a large multilingual ST translation corpus available . It contains translations from English into fourteen additional languages and is compiled from TED Talks. mTEDx (Salesky et al., 2021) is one such multilingual dataset from TED talks.
@@
    Kosp2e (Cho et al., 2021) is a Korean (ko) to English(en) ST translation corpus, which contains Korean speech with parallel English texts. The corpus contains data from four different domains: Zeroth from news/newspaper, KSS (Park, 2018) from textbooks, emphStyleKQC (Cho et al., 2022) from AI applications, and Covid-ED (Lee et al., 2021) from covid diaries of people which have emotions.
@@
    BSTC (Zhang et al., 2021) is a Baidu Speech Translation Corpus, a large-scale Chinese-English speech translation dataset. This dataset is constructed based on a collection of licensed videos of talks or lectures, their manual transcripts, and translations into English, as well as automated transcripts by an automatic speech recognition (ASR) model.
  • 9.
@@
  • 11.
    FLEURS (Conneau et al., 2022) FLEURS stands as a multilingual speech dataset, offering parallel recordings across 102 languages. Developed as an extension of the FLoRes-101 MT benchmark, it encompasses about 12 hours of annotated speech data for each language.
  • 12.
    Indic-TEDST Sethiya et al. (2024) is a low-resource ST translation dataset across 9 Indic languages: Bengali (bn), Gujarati (gu), Hindi (hi), Kannada (kn), Malayalam (ml), Marathi (mr), Punjabi (pa), Tamil (ta), and Telugu (te).
Besides these popular ST datasets, there are some other smaller size datasets such as Fisher(Cieri et al., 2004), Call-Home [^13] https://ca.talkbank.org/access/CallHome/eng.html, Gordard Corpus(Godard et al., 2018), Glosse Audio Corpus [^14] https://ainu.ninjal.ac.jp/folklore/en/, BTEC  [^15] http://universal.elra.info/product_info.php?cPath=37_39&products_id=80, WSJ [^16] https://catalog.ldc.upenn.edu/LDC93s6a, IWSLT [^17] https://iwslt.org/, Miami Corpus(Deuchar, 2008), and MSLT Corpus (Federmann and Lewis, 2016).
9.2 Toolkits for ST
To facilitate building and training ST models, various researchers have proposed a few toolkits. The toolkits for ST create an environment where the dataset for ST tasks can be pre-processed, and models can be trained, fine-tuned, and evaluated. We provide a short description of these toolkits to make the survey a place for a one-stop-shop for ST modeling.
@@ 9.2 Toolkits for ST
    EspNet-ST [^19] https://github.com/espnet/espnet toolkit (Inaguma et al., 2020b) is developed as there was no toolkit available for performing the sub-tasks of ST. EspNet-ST provides ASR, LM, E2E-ST, Cascade-ST, MT, and TTS along with examples. It also provided pre-trained transformer-based models on various datasets like MUST-C, Libri-trans, Fisher, CALL-HOME, and How2.
  • 3.